### 25 Mar 2026 v1

@@ 3.2 Cell Clustering
Fig.˜5a represents ARI of cell clustering based on the imputed and ground truth data for the 15 imputation methods in terms of 26 real and 4 simulated datasets. Particularly, it demonstrates the consistency between cell clustering using imputed and ground truth data. $\mathrm{ARI}=1$ shows clusters match perfectly, $\mathrm{ARI}=0$ shows cell clustering performance is equivalent to randomly assigning clusters, and $\mathrm{ARI}<0$ shows cell clustering performance is worse than randomly assigning clusters. Out of 15 distinct imputation methods, scLRTC exhibits the highest ARI scores in 13 datasets. In addition, 12 methods, namely PbImpute, scImpute, AcImpute, scTsI, MAGIC, WEDGE, scIDPMs, scMultiGAN, scIGANs, scGNN, CPARI, and Bubble, show moderate ARI scores. On the other hand, PBLR and stDiff show the lowest ARI scores for 8 datasets. Moreover, for 12 datasets, including sc_10x_5cl, hca_10x_tissue, cellmix1, sc_celseq2, hcc, petropoulos, chu_time_course, chen, romanov, usokin, zeisel, and baron, none of the 15 methods exceed the ARI scores of the masked baseline. This indicates that imputation does not necessarily improve cell clustering and can even degrade performance compared to using the masked baseline data.
Fig.˜5b represents SC for the 15 imputation methods in terms of 26 real and 4 simulated datasets. $\mathrm{SC}=1$ represents cells are perfectly assigned to highly dense and isolated clusters, $\mathrm{SC}=0$ represents cells are located at the boundaries between 2 clusters, and $\mathrm{SC}<0$ represents cells are inaccurately assigned to clusters. The comparison of SC scores among the 15 methods shows that MAGIC and WEDGE achieve the best performance for 9 datasets. On the other hand, PBLR shows the worst SC scores for 13 datasets. The remaining 12 methods, namely PbImpute, scImpute, AcImpute, scTsI, scLRTC, scIDPMs, stDiff, scMultiGAN, scIGANs, scGNN, CPARI, and Bubble, show moderate SC scores.
@@
Figs.˜6 and 7 represent UMAP visualization of the cell clustering for the 15 imputation methods in terms of 4 real datasets, and 4 simulated datasets, respectively. The datasets with the highest number of cells are selected from 4 protocols, including 10x Chromium, CEL-seq2, SMART-seq2, and Drop-seq, to perform a qualitative evaluation. A qualitative assessment of UMAP plots provides a visual perspective on cluster consistency and coherency that complements the quantitative evaluations. The comparison of cluster structures among the 15 methods shows that MAGIC and WEDGE produce the most visually coherent clusters across 4 real datasets, including sc_celseq2_5cl_p1 with small samples. On the other hand, stDiff and PBLR show the least coherent cluster structures across 4 real datasets, with clusters appearing fragmented or poorly separated compared to the cell clustering based on the ground truth data. In the simulated datasets, 5 methods, namely PbImpute, scImpute, MAGIC, WEDGE, and scMultiGAN, maintain visually distinct clusters across 4 simulated datasets. On the other hand, the remaining 10 methods show limited ability to recover the cluster structure of the ground truth data. This suggests that these 10 methods have lower robustness to dropout-induced sparsity.
Tables˜3 and 4 report NMI and purity scores, respectively, comparing cell clustering results from the imputed data with those from ground truth data across the 15 imputation methods for the 26 real and 4 simulated datasets. An NMI score of $1$ indicates perfect agreement between the 2 clustering results, whereas 0 indicates independence. Similarly, a purity score of $1$ indicates that each predicted cluster contains cells from a single ground truth cluster, while 0 indicates complete mixing of cells from different ground truth clusters. The analyses of NMI and purity scores among the 15 imputation methods show that these scores are largely consistent with the ARI scores, where scLRTC exhibits the best performance in 13 datasets, PBLR and stDiff show the lowest NMI and purity scores in 8 datasets, and the remaining 12 methods, including PbImpute, scImpute, AcImpute, scTsI, MAGIC, WEDGE, scIDPMs, scMultiGAN, scIGANs, scGNN, CPARI, and Bubble, show moderate NMI and purity scores. This consistency across multiple cell clustering metrics reinforces the robustness of the observed performance differences among the 15 methods.
@@ 3.3 DE analysis
Following Hou et al. [19], 3 different DE analyses are conducted, namely DE enrichment analysis, null DE analysis, and effect size analysis, as illustrated in Fig.˜8. The DE enrichment analysis evaluates how well DEGs from imputed scRNA-seq data recover DEGs identified from bulk RNA-seq data, which serve as the ground truth DEGs. The DEGs from imputed scRNA-seq data are ranked by $p$ values or LFC if there is a tie for $p$ values, and the IoU between the bulk RNA-seq DEGs and the top $10i$ imputed DEGs is computed for $i$ from 1 to 100. The average IoU across all 100 values of $i$ is used to measure the performance. The null DE analysis assesses the robustness of imputed data to false positive DEGs. Ideally, DEGs should not be identified when control and target groups belong to the same cell population, and any identified DEGs can be treated as false positive DEGs under such conditions. The imputed scRNA-seq data is filtered to a single cell type or cell line to obtain a homogeneous cell population. From this cell population, $N_{1}$ and $N_{2}$ cells are randomly sampled as control and target groups, respectively, where $N_{1}\leq N_{2}$ and $N_{1},N_{2}\in\{10,50,100\}$. All 6 combinations, namely $(N_{1},N_{2})=(10,10),(10,50),(10,100),(50,50),(50,100),(100,100)$, are tested to assess robustness across varying sample sizes and group balances, and DE analysis is performed between the control and target groups. The number of false positive DEGs is used to measure the performance, where a lower count indicates greater robustness. The effect size analysis evaluates whether DEGs identified from imputed scRNA-seq data capture genes with both high and low LFC in bulk RNA-seq data. Here, LFC is defined as $\mathrm{LFC}=\log_{2}(y_{\mathrm{target}}/y_{\mathrm{control}})$, where $y_{\mathrm{control}}$ and $y_{\mathrm{target}}$ represent the gene expression values of the control and target groups, respectively. Genes in the upper and lower $10\,\%$ of the LFC distribution from bulk RNA-seq data are defined as high- and low-LFC genes, respectively. As in the DE enrichment analysis, the IoU is used to measure the overlap between high- or low-LFC genes from bulk RNA-seq data and the DEGs from imputed scRNA-seq data.
@@ 3.4 Marker Gene Analysis
Fig.˜10 shows marker gene expression of 4 cell types, namely T cell, B cell, natural killer (NK) cell, and monocyte, in hca_10x_tissue for 15 distinct methods. Generally, CD3D and CD3E are considered to be marker genes for T cells, CD79A and MS4A1 for B cells, NKG7 and GNLY for NK cells, and CD14 and LYZ for monocytes [17, 118]. The analysis of marker gene expression reveals that scImpute, MAGIC, scIGANs, and CPARI achieve the best performance, as these methods show strong expression levels of marker genes in the corresponding cell types. In addition, 6 methods, namely PbImpute, AcImpute, scTsI, WEDGE, scMultiGAN, and Bubble, show moderate performance, as these methods show relatively strong expression levels of marker genes for B cells and monocytes. Conversely, 5 methods, namely PBLR, scLRTC, scIDPMs, stDiff, and scGNN, show the worst performance, as they produce similar marker gene expression levels across the 4 cell types or show near-zero expression levels.
@@ 3.5 Trajectory Analysis
Figs. 12a and b show POS and KRCC for 15 imputation methods in terms of 2 datasets, namely petropoulos and chu_time_course. High POS and KRCC represent proximity to true cell development labels. In each plot, points close to the dashed line represent consistent performance across the 2 datasets. The comparison of POS and KRCC across 15 methods evaluated on 2 datasets shows that PbImpute, scImpute, scLRTC, CPARI, and Bubble achieve relatively high performance in both datasets. In addition, 7 methods, namely AcImpute, scTsI, MAGIC, PBLR, WEDGE, scMultiGAN, and scGNN, show moderate performance. Conversely, scIDPMs and scIGANs show the worst performance in petropoulos dataset, while stDiff shows the worst performance in chu_time_course dataset. However, the masked baseline shows relatively high performance, and 10 methods, including AcImpute, scTsI, MAGIC, PBLR, WEDGE, scIDPMs, stDiff, scMultiGAN, scIGANs, and scGNN, do not exceed it.
Fig.˜12c shows UMAP visualizations of trajectory analysis for 15 imputation methods in terms of 2 datasets. The qualitative analyses through the UMAP visualizations show that the pseudotime changes gradually for PbImpute and scImpute, with cells colored by pseudotime progressing smoothly. On the other hand, the UMAP visualizations of AcImpute, scTsI, PBLR, WEDGE, scIDPMs, stDiff, scMultiGAN, scIGANs, and CPARI show that the pseudotime does not change smoothly, with cells of different time points appearing intermixed along the trajectory. However, since trajectory analysis is performed in higher dimensions than UMAP, POS and KRCC results may not be fully reflected in the 2-dimensional UMAP projections.
@@ 3.6 Cell Type Annotation
Fig.˜13 shows ACC, PR, RC, and F1 of cell type annotation for the 15 imputation methods in terms of 6 datasets, namely sc_10x_5cl, hca_10x_tissue, chu_cell_type, baron, encode_fluidigm_5cl, and bladder. 1D-CNN and scGPT [109] are used to annotate cell types. Higher ACC represents better overall accuracy of cell type annotations, higher PR represents better precision in identifying true cell types, higher RC represents better recall of true cell types, and higher F1 represents a better balance between precision and recall.
@@ 4 Discussion
Table˜5 summarizes the performance of the 15 imputation methods across 6 downstream tasks. MAGIC shows the best overall performance with the highest performance in 3 tasks, namely cell clustering, marker gene analysis, and cell type annotation, and moderate performance in the remaining 3 tasks. In addition, scImpute and WEDGE show moderately high overall performance with the best performance in 2 tasks, and moderate performance in the remaining 4 tasks. Similarly, PbImpute, AcImpute, scTsI, scMultiGAN, CPARI, and Bubble show moderately low overall performance, with the best performance in 0 or 1 tasks, and moderate performance in the remaining 5 or 6 tasks. Conversely, PBLR, scLRTC, scIDPMs, stDiff, scIGANs, and scGNN show the worst overall performance, with the worst performance in 1 to 3 tasks. We find that traditional methods, such as scImpute, MAGIC, and WEDGE, show the best or moderately high overall performance across the 6 tasks, whereas none of the 7 DL-based methods reach the same level of overall performance. This suggests that traditional methods may be more effective at preserving biologically meaningful information across a wide range of downstream analyses, while the performance of DL-based methods can be more variable and may require careful tuning and validation for specific tasks.
@@
In numerical gene expression recovery, the performance of imputation methods varies significantly across methods. scTsI, PBLR, and WEDGE achieve the best overall LND performance, with medians consistently close to 0. scTsI achieves this through its two-stage strategy, which first imputes dropouts using $k$-NN-based averaging across neighboring cells and genes, and refines the initial estimates through ridge regression. PBLR preserves the data structure through its cell sub-population-based bounded low-rank matrix recovery. The boundaries constrain reconstructed values within biologically plausible ranges. WEDGE shows superior LND performance likely due to its biased low-rank matrix-based approach that assigns low weights to 0 elements and minimizes approximation error for nonzero elements. This approach effectively separates true biological signal from dropouts without over-imputing zero entries. scTsI and WEDGE also exhibit the lowest MAE and MedAE, which further supports this finding. Protocol-wise analysis reveals that WEDGE maintains LND values closest to zero with compact distributions across all protocols, which indicates that it is the most protocol-robust approach among the 15 methods. For comparison with bulk RNA-seq data, WEDGE also achieves the best overall performance, while MAGIC shows the highest cell line-level correlation through its diffusion-based smoothing, which enforces locally coherent expression profiles aligned with averaged bulk patterns. In contrast, scMultiGAN exhibits the worst correlation with bulk RNA-seq data at both the pseudo-bulk and cell line levels despite its moderate numerical recovery performance from ground truth data. This highlights a trade-off between numerical accuracy and biological fidelity. Its dual-GAN architecture, which minimizes numerical recovery error, may overfit to ground truth distributions at the expense of generalization to bulk RNA-seq data. The over-imputation by scIDPMs may arise from its iterative denoising process that pushes sparse dropout entries toward higher-density non-zero modes across multiple sequential denoising steps. scLRTC consistently under-imputes because its low-rank matrix-based approach compresses the dynamic range of highly variable genes. scIGANs exhibits the highest protocol-wise MAE and MedAE that significantly exceed the masked baseline, and also shows the worst correlation with bulk RNA-seq data. This worst performance likely arises because its adversarially trained generator learns conservative mappings that fail to generalize across protocols and sequencing modalities. Furthermore, methods show largely consistent behavior on 10x Chromium, CEL-seq2, and Drop-seq datasets, while SMART-seq, SMART-seq2, and Fluidigm C1 datasets demonstrate higher instability. This is likely because the latter 3 protocols use only read-counts without UMIs. The absence of UMIs leads to increased technical noise that makes accurate imputation more challenging [17, 119].
@@
In trajectory analysis, we evaluate the ability of imputation methods to preserve the temporal ordering of cells by comparing inferred pseudotime from imputed data with true cellular development time labels. The results reveal that PbImpute, scImpute, scLRTC, CPARI, and Bubble consistently preserve the temporal ordering of cells across the 2 datasets. This shows that 2 model-based methods, 1 low-rank matrix-based method, and 2 AE-based methods perform well in trajectory analysis. This may be due to the fact that these methods preserve the relative expression differences along developmental trajectories, which is critical for accurate pseudotime inference. The model-based methods, scImpute and PbImpute, selectively target dropout events while preserving the original expression values, which helps maintain the gradual expression changes that define developmental trajectories. scLRTC reconstructs gene expression values through its low-rank tensor completion that captures global gene expression patterns while preserving the distinct expression signatures that define developmental trajectories. The AE-based methods, CPARI and Bubble, learn latent representations using autoencoder architectures that retain the continuous structure of developmental progression without collapsing intermediate states. Conversely, scIDPMs, stDiff, and scIGANs show the worst performance. The diffusion-based methods, namely scIDPMs and stDiff, use an iterative denoising process trained to denoise expression values toward high-density modes, which may collapse the subtle expression gradients between adjacent developmental stages into discrete expression states. This disrupts the continuous pseudotime ordering that trajectory inference relies on, as cells at intermediate developmental stages are pushed toward the expression profiles of more mature or earlier stages. For scIGANs, its adversarial training aims to match the overall data distribution, but the GAN-based generation process may suffer from mode collapse that concentrates imputed values around high-density expression states rather than preserving the continuous gradients between developmental stages. This can disrupt the temporal ordering of cells by homogenizing expression profiles at intermediate stages. This is consistent with the UMAP visualizations, where cells of different time points appear intermixed along the trajectory for these methods. Furthermore, 10 methods, including AcImpute, scTsI, MAGIC, PBLR, WEDGE, scIDPMs, stDiff, scMultiGAN, scIGANs, and scGNN, fail to outperform the masked baseline, which highlights that imputation can distort the underlying developmental structure of scRNA-seq data and lead to less accurate trajectory inference than simply using the original data. This suggests that trajectory analysis, which depends on the preservation of continuous expression gradients rather than discrete cluster boundaries, is particularly sensitive to imputation artifacts that alter the relative ordering of expression values along developmental axes.

### 13 Apr 2026 v2

@@ 2.2 Cell Clustering
Fig.˜3a represents adjusted rand index (ARI) of cell clustering based on the imputed and ground truth data for the 15 imputation methods in terms of 26 real and 4 simulated datasets. Particularly, it demonstrates the consistency between cell clustering using imputed and ground truth data. $\mathrm{ARI}=1$ shows clusters match perfectly, $\mathrm{ARI}=0$ shows cell clustering performance is equivalent to randomly assigning clusters, and $\mathrm{ARI}<0$ shows cell clustering performance is worse than randomly assigning clusters. Out of 15 distinct imputation methods, scLRTC exhibits the highest ARI scores in 13 datasets. In addition, 12 methods, namely PbImpute, scImpute, AcImpute, scTsI, MAGIC, WEDGE, scIDPMs, scMultiGAN, scIGANs, scGNN, CPARI, and Bubble, show moderate ARI scores. On the other hand, PBLR and stDiff show the lowest ARI scores for 8 datasets. Moreover, for 12 datasets, including sc_10x_5cl, hca_10x_tissue, cellmix1, sc_celseq2, hcc, petropoulos, chu_time_course, chen, romanov, usokin, zeisel, and baron, none of the 15 methods exceed the ARI scores of the masked baseline. This indicates that imputation does not necessarily improve cell clustering and can even degrade performance compared to using the masked baseline data.
Fig.˜3b represents silhouette coefficient (SC) for the 15 imputation methods in terms of 26 real and 4 simulated datasets. $\mathrm{SC}=1$ represents cells are perfectly assigned to highly dense and isolated clusters, $\mathrm{SC}=0$ represents cells are located at the boundaries between 2 clusters, and $\mathrm{SC}<0$ represents cells are inaccurately assigned to clusters. The comparison of SC scores among the 15 methods shows that MAGIC and WEDGE achieve the best performance for 9 datasets. On the other hand, PBLR shows the worst SC scores for 13 datasets. The remaining 12 methods, namely PbImpute, scImpute, AcImpute, scTsI, scLRTC, scIDPMs, stDiff, scMultiGAN, scIGANs, scGNN, CPARI, and Bubble, show moderate SC scores.
@@
Figs.˜4 and 5 represent uniform manifold approximation and projection (UMAP) visualization of the cell clustering for the 15 imputation methods in terms of 4 real datasets, and 4 simulated datasets, respectively. The datasets with the highest number of cells are selected from 4 protocols, including 10x Chromium, CEL-seq2, SMART-seq2, and Drop-seq, to perform a qualitative evaluation. A qualitative assessment of UMAP plots provides a visual perspective on cluster consistency and coherency that complements the quantitative evaluations. The comparison of cluster structures among the 15 methods shows that MAGIC and WEDGE produce the most visually coherent clusters across 4 real datasets, including sc_celseq2_5cl_p1 with small samples. On the other hand, stDiff and PBLR show the least coherent cluster structures across 4 real datasets, with clusters appearing fragmented or poorly separated compared to the cell clustering based on the ground truth data. In the simulated datasets, 5 methods, namely PbImpute, scImpute, MAGIC, WEDGE, and scMultiGAN, maintain visually distinct clusters across 4 simulated datasets. On the other hand, the remaining 10 methods show limited ability to recover the cluster structure of the ground truth data. This suggests that these 10 methods have lower robustness to dropout-induced sparsity.
@@
Tables˜2 and 3 report normalized mutual information (NMI) and purity scores, respectively, comparing cell clustering results from the imputed data with those from ground truth data across the 15 imputation methods for the 26 real and 4 simulated datasets. An NMI score of $1$ indicates perfect agreement between the 2 clustering results, whereas 0 indicates independence. Similarly, a purity score of $1$ indicates that each predicted cluster contains cells from a single ground truth cluster, while 0 indicates complete mixing of cells from different ground truth clusters. The analyses of NMI and purity scores among the 15 imputation methods show that these scores are largely consistent with the ARI scores, where scLRTC exhibits the best performance in 13 datasets, PBLR and stDiff show the lowest NMI and purity scores in 8 datasets, and the remaining 12 methods, including PbImpute, scImpute, AcImpute, scTsI, MAGIC, WEDGE, scIDPMs, scMultiGAN, scIGANs, scGNN, CPARI, and Bubble, show moderate NMI and purity scores. This consistency across multiple cell clustering metrics reinforces the robustness of the observed performance differences among the 15 methods.
@@ 2.3 DE Analysis
Following Hou et al. [19], 3 different DE analyses are conducted, namely DE enrichment analysis, null DE analysis, and effect size analysis, as illustrated in Fig.˜6. The DE enrichment analysis evaluates how well differentially expressed genes (DEGs) from imputed scRNA-seq data recover DEGs identified from bulk RNA-seq data, which serve as the ground truth DEGs. The DEGs from imputed scRNA-seq data are ranked by $p$ values or log-fold change (LFC) if there is a tie for $p$ values, and the intersection over union (IoU) between the bulk RNA-seq DEGs and the top $10i$ imputed DEGs is computed for $i$ from 1 to 100. The average IoU across all 100 values of $i$ is used to measure the performance. The null DE analysis assesses the robustness of imputed data to false positive DEGs. Ideally, DEGs should not be identified when control and target groups belong to the same cell population, and any identified DEGs can be treated as false positive DEGs under such conditions. The imputed scRNA-seq data is filtered to a single cell type or cell line to obtain a homogeneous cell population. From this cell population, $N_{1}$ and $N_{2}$ cells are randomly sampled as control and target groups, respectively, where $N_{1}\leq N_{2}$ and $N_{1},N_{2}\in\{10,50,100\}$. All 6 combinations, namely $(N_{1},N_{2})=(10,10),(10,50),(10,100),(50,50),(50,100),(100,100)$, are tested to assess robustness across varying sample sizes and group balances, and DE analysis is performed between the control and target groups. The number of false positive DEGs is used to measure the performance, where a lower count indicates greater robustness. The effect size analysis evaluates whether DEGs identified from imputed scRNA-seq data capture genes with both high and low LFC in bulk RNA-seq data. Here, LFC is defined as $\mathrm{LFC}=\log_{2}(y_{\mathrm{target}}/y_{\mathrm{control}})$, where $y_{\mathrm{control}}$ and $y_{\mathrm{target}}$ represent the gene expression values of the control and target groups, respectively. Genes in the upper and lower $10\,\%$ of the LFC distribution from bulk RNA-seq data are defined as high- and low-LFC genes, respectively. As in the DE enrichment analysis, the IoU is used to measure the overlap between high- or low-LFC genes from bulk RNA-seq data and the DEGs from imputed scRNA-seq data.
@@ 2.4 Marker Gene Analysis
Fig.˜8 shows marker gene expression of 4 cell types, namely T cell, B cell, natural killer (NK) cell, and monocyte, in hca_10x_tissue for 15 distinct methods. Generally, CD3D and CD3E are considered to be marker genes for T cells, CD79A and MS4A1 for B cells, NKG7 and GNLY for NK cells, and CD14 and LYZ for monocytes [17, 60]. The analysis of marker gene expression reveals that scImpute, MAGIC, scIGANs, and CPARI achieve the best performance, as these methods show strong expression levels of marker genes in the corresponding cell types. In addition, 6 methods, namely PbImpute, AcImpute, scTsI, WEDGE, scMultiGAN, and Bubble, show moderate performance, as these methods show relatively strong expression levels of marker genes for B cells and monocytes. Conversely, 5 methods, namely PBLR, scLRTC, scIDPMs, stDiff, and scGNN, show the worst performance, as they produce similar marker gene expression levels across the 4 cell types or show near-zero expression levels.
@@ 2.5 Trajectory Analysis
Figs. 10a and b show pseudo-temporal ordering score (POS) and Kendall’s rank correlation coefficient (KRCC) for 15 imputation methods in terms of 2 datasets, namely petropoulos and chu_time_course. High POS and KRCC represent proximity to true cell development labels. The comparison of POS and KRCC across 15 methods evaluated on 2 datasets shows that PbImpute, scImpute, scLRTC, CPARI, and Bubble achieve relatively high performance in both datasets. In addition, 7 methods, namely AcImpute, scTsI, MAGIC, PBLR, WEDGE, scMultiGAN, and scGNN, show moderate performance. Conversely, scIDPMs and scIGANs show the worst performance in petropoulos dataset, while stDiff shows the worst performance in chu_time_course dataset. However, the masked baseline shows relatively high performance, and 10 methods, including AcImpute, scTsI, MAGIC, PBLR, WEDGE, scIDPMs, stDiff, scMultiGAN, scIGANs, and scGNN, do not exceed it.
Fig.˜10c shows UMAP visualizations of trajectory analysis for 15 imputation methods in terms of 2 datasets. The qualitative analyses through the UMAP visualizations show that the pseudotime changes gradually for PbImpute and scImpute, with cells colored by pseudotime progressing smoothly. On the other hand, the UMAP visualizations of AcImpute, scTsI, PBLR, WEDGE, scIDPMs, stDiff, scMultiGAN, scIGANs, and CPARI show that the pseudotime does not change smoothly, with cells of different time points appearing intermixed along the trajectory. However, since trajectory analysis is performed in higher dimensions than UMAP, POS and KRCC results may not be fully reflected in the 2-dimensional UMAP projections.
@@ 2.6 Cell Type Annotation
Fig.˜11 shows the macro accuracy (ACC), macro precision (PR), macro recall (RC), and macro F1 score (F1) of cell type annotation for the 15 imputation methods in terms of 6 datasets, namely sc_10x_5cl, hca_10x_tissue, chu_cell_type, baron, encode_fluidigm_5cl, and bladder, using 1D convolutional neural network (1D-CNN) and scGPT [61] for annotation. Higher ACC represents better overall accuracy of cell type annotations, higher PR represents better precision in identifying true cell types, higher RC represents better recall of true cell types, and higher F1 represents a better balance between PR and RC.
@@ 3 Discussion
Table˜4 summarizes the performance of the 15 imputation methods across 6 downstream tasks. MAGIC shows the best overall performance with the highest performance in 3 tasks, namely cell clustering, marker gene analysis, and cell type annotation, and moderate performance in the remaining 3 tasks. In addition, scImpute and WEDGE show moderately high overall performance with the best performance in 2 tasks, and moderate performance in the remaining 4 tasks. Similarly, PbImpute, AcImpute, scTsI, scMultiGAN, CPARI, and Bubble show moderately low overall performance, with the best performance in 0 or 1 tasks, and moderate performance in the remaining 5 or 6 tasks. Conversely, PBLR, scLRTC, scIDPMs, stDiff, scIGANs, and scGNN show the worst overall performance, with the worst performance in 1 to 3 tasks. We find that traditional methods, such as scImpute, MAGIC, and WEDGE, show the best or moderately high overall performance across the 6 tasks, whereas none of the 7 DL-based methods reach the same level of overall performance. This suggests that traditional methods may be more effective at preserving biologically meaningful information across a wide range of downstream analyses, while the performance of DL-based methods can be more variable and may require careful tuning and validation for specific tasks.
@@
In numerical gene expression recovery, the performance of imputation methods varies significantly across methods. scTsI, PBLR, and WEDGE achieve the best overall LND performance, with medians consistently close to 0. scTsI achieves this through its two-stage strategy, which first imputes dropouts using $k$-nearest neighbor ($k$-NN)-based averaging across neighboring cells and genes, and refines the initial estimates through ridge regression. PBLR preserves the data structure through its cell sub-population-based bounded low-rank matrix recovery. The boundaries constrain reconstructed values within biologically plausible ranges. WEDGE shows superior LND performance likely due to its biased low-rank matrix-based approach that assigns low weights to 0 elements and minimizes approximation error for nonzero elements. This approach effectively separates true biological signal from dropouts without over-imputing zero entries. scTsI and WEDGE also exhibit the lowest MAE and MedAE, which further supports this finding. Protocol-wise analysis reveals that WEDGE maintains LND values closest to zero with compact distributions across all protocols, which indicates that it is the most protocol-robust approach among the 15 methods. For comparison with bulk RNA-seq data, WEDGE also achieves the best overall performance, while MAGIC shows the highest cell line-level correlation through its diffusion-based smoothing, which enforces locally coherent expression profiles aligned with averaged bulk patterns. In contrast, scMultiGAN exhibits the worst correlation with bulk RNA-seq data at both the pseudo-bulk and cell line levels despite its moderate numerical recovery performance from ground truth data. This highlights a trade-off between numerical accuracy and biological fidelity. Its dual-GAN architecture, which minimizes numerical recovery error, may overfit to ground truth distributions at the expense of generalization to bulk RNA-seq data. The over-imputation by scIDPMs may arise from its iterative denoising process that pushes sparse dropout entries toward higher-density non-zero modes across multiple sequential denoising steps. scLRTC consistently under-imputes because its low-rank matrix-based approach compresses the dynamic range of highly variable genes. scIGANs exhibits the highest protocol-wise MAE and MedAE that significantly exceed the masked baseline, and also shows the worst correlation with bulk RNA-seq data. This worst performance likely arises because its adversarially trained generator learns conservative mappings that fail to generalize across protocols and sequencing modalities. Furthermore, methods show largely consistent behavior on 10x Chromium, CEL-seq2, and Drop-seq datasets, while SMART-seq, SMART-seq2, and Fluidigm C1 datasets demonstrate higher instability. This is likely because the latter 3 protocols use only read-counts without UMIs. The absence of UMIs leads to increased technical noise that makes accurate imputation more challenging [17, 62].
@@
In trajectory analysis, we evaluate the ability of imputation methods to preserve the temporal ordering of cells by comparing inferred pseudotime from imputed data with true cellular development time labels. The results reveal that PbImpute, scImpute, scLRTC, CPARI, and Bubble consistently preserve the temporal ordering of cells across the 2 datasets. This shows that 2 model-based methods, 1 low-rank matrix-based method, and 2 AE-based methods perform well in trajectory analysis. This may be due to the fact that these methods preserve the relative expression differences along developmental trajectories, which is critical for accurate pseudotime inference. The model-based methods, scImpute and PbImpute, selectively target dropout events while preserving the original expression values, which helps maintain the gradual expression changes that define developmental trajectories. scLRTC reconstructs gene expression values through its low-rank tensor completion that captures global gene expression patterns while preserving the distinct expression signatures that define developmental trajectories. The AE-based methods, CPARI and Bubble, learn latent representations using AE architectures that retain the continuous structure of developmental progression without collapsing intermediate states. Conversely, scIDPMs, stDiff, and scIGANs show the worst performance. The diffusion-based methods, namely scIDPMs and stDiff, use an iterative denoising process trained to denoise expression values toward high-density modes, which may collapse the subtle expression gradients between adjacent developmental stages into discrete expression states. This disrupts the continuous pseudotime ordering that trajectory inference relies on, as cells at intermediate developmental stages are pushed toward the expression profiles of more mature or earlier stages. For scIGANs, its adversarial training aims to match the overall data distribution, but the GAN-based generation process may suffer from mode collapse that concentrates imputed values around high-density expression states rather than preserving the continuous gradients between developmental stages. This can disrupt the temporal ordering of cells by homogenizing expression profiles at intermediate stages. This is consistent with the UMAP visualizations, where cells of different time points appear intermixed along the trajectory for these methods. Furthermore, 10 methods, including AcImpute, scTsI, MAGIC, PBLR, WEDGE, scIDPMs, stDiff, scMultiGAN, scIGANs, and scGNN, fail to outperform the masked baseline, which highlights that imputation can distort the underlying developmental structure of scRNA-seq data and lead to less accurate trajectory inference than simply using the original data. This suggests that trajectory analysis, which depends on the preservation of continuous expression gradients rather than discrete cluster boundaries, is particularly sensitive to imputation artifacts that alter the relative ordering of expression values along developmental axes.